Shift report 20011123 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Fri Nov 23 2001 - 15:47:10 EST

  • Next message: Apache: "Shift report 20011124 00:00-08:00"

    ------------------------------------------------------------
    Date: Fri Nov 23 15:47:09 2001, Shift: 08:00-16:00
    Supervisor: rw, with: 
    Log book#11 pages 28 - 28
    
    Shift Summary:
    ------------------------------------------------------------
    1. Beam from the previous shift was dumped at 12:25 after 1.5h
    beam study. Yeloow ring decayed much faster than the Blue ring.
    At 13:15 beam came back with Blue/Yellow=24/37 ZDC rate of 500.
    ZDC rate increased to 800 after steering but one hour later.
    Yellow ring behaves much better than before.
    
    2.	run 	trigger 6   	setup
    	5956	105K		same as the previous run
    				(FS 12deg.-2/3A, MRS 90deg.-6B) 
    	5957	129K		same as above
    	5958	 28K		during Beam study
    	5959    	TPC pedestal run
    	5960		DVM
    	5961		Pedestal run
    	5962     64K		same as before
    	5963	 82K		same as above
    	5964	running
    3. No new detector problems have been detected.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Fri Nov 23 2001 - 15:48:15 EST